# Predict-and-Optimize Robust Unit Commitment with Statistical Guarantees via Weight Combination

### v2

@@ Abstract
The growing uncertainty from renewable power and electricity demand brings significant challenges to unit commitment (UC). While various advanced forecasting and optimization methods have been developed to predict better and address this uncertainty, most previous studies treat forecasting and optimization as separate tasks. This separation can lead to suboptimal results due to misalignment between the objectives of the two tasks. To overcome this challenge, we propose a robust UC framework that integrates forecasting and optimization processes while ensuring statistical guarantees. In the forecasting stage, we combine multiple predictions derived from diverse data sources and methodologies for an improved prediction, aiming to optimize the UC performance. In the optimization stage, the combined prediction is used to construct an uncertainty set with statistical guarantees, based on which the robust UC model is formulated. The optimal robust UC solution provides feedback to refine the weight used for combining multiple predictions. To solve the proposed integrated forecasting-optimization framework efficiently and effectively, we develop a neural network-based surrogate model for acceleration and introduce a reshaping method for the uncertainty set based on the optimization result to reduce conservativeness. Case studies on modified IEEE 30-bus and 118-bus systems demonstrate the advantages of the proposed approach.
Index Terms:
unit commitment, data-driven robust optimization, statistical guarantee, predict-and-optimize, surrogate model
@@ I Introduction
If given the probability distribution of uncertainty, stochastic programming (SP) can be applied to UC problems to determine the optimal strategy [1]. However, an accurate probability distribution is difficult to obtain in reality and the inaccuracy of the distribution may jeopardize the feasibility and optimality of SP solutions. Robust optimization (RO) handles this problem by focusing on the worst-case scenario in a pre-defined uncertainty set. Fruitful research has been done in this area: A multi-stage robust UC model was proposed and solved by robust dual dynamic programming in [2]. Polyhedra were used to replace rectangular uncertainty sets to model the uncertainty correlations in [3]. However, RO can be overly conservative when the uncertainty set contains all the possible uncertainty realizations, because extreme scenarios rarely happen but can cause a steep cost increase. The method in [4] estimated quantiles from historical data to establish rectangular uncertainty sets, in which extreme scenarios were excluded. In [5], uncertainty budgets were used to restrict the uncertainty set. A decision-dependent uncertainty set was constructed in [6] to consider the impact of pricing strategies on demand. However, estimated quantiles and uncertainty budgets may also be inaccurate, and the resulting uncertainty set lacks a statistical guarantee. This makes the obtained robust UC strategy less reliable and trustworthy.
@@ II Integrated Forecasting and Optimization Framework
In this section, we first propose an integrated forecasting and optimization framework for a two-stage decision-making problem under uncertainty. In fact, the proposed framework is generally applicable and not restricted to UC. Then in Section III, we introduce the detailed UC model and explain how it fits within the proposed framework.
@@ II-A Forecasting
Consider the forecasting task for a multivariate stochastic process $\mathbf{u}:=(\mathbf{u}_{1},\mathbf{u}_{2},\dots,\mathbf{u}_{t},\dots)$, where $\mathbf{u}_{t}=(u_{it};i\in\mathcal{I})$ is the vector of uncertain variables at time $t$, and $\mathcal{I}$ is the index set. A forecast is then a collection of future values over the horizon of interest, denoted by $\hat{\mathbf{U}}:=(\hat{\mathbf{u}}_{1},\hat{\mathbf{u}}_{2},\dots,\hat{\mathbf{u}}_{T})$. The hat notation indicates these are forecasts, and hence estimates of such future values. Additionally, the corresponding ground truth is denoted by $\mathbf{U}$.
Multiple forecasting methods can be used simultaneously. The prediction by method $m\in\mathcal{M}$ is denoted by $\hat{\mathbf{U}}^{(m)}$, and $\mathcal{M}$ is the set of methods. The final prediction is a convex combination of these predictions, that is, $\hat{\mathbf{U}}(w)=\sum_{m\in\mathcal{M}}w^{(m)}\hat{\mathbf{U}}^{(m)}$, where $w=(w^{(m)};m\in\mathcal{M})$ is the weight vector. It satisfies $w\in\mathcal{W}:=\{w~|~w\geq 0,\sum_{m\in\mathcal{M}}w^{(m)}=1\}$. Using the ground truth $\mathbf{U}$, the forecast error is calculated as $\mathbf{E}(w)=\mathbf{U}-\hat{\mathbf{U}}(w)$, which also depends on the weight $w$. Forecasting methods are not the focus of this paper, and we apply multiple forecasting methods in case studies based on the related literature.
@@ II-D2 MILP-Based Weight Optimization
where $z$ is an auxiliary binary variable and $M>0$ is a sufficiently large constant. When $z=0$, (4a) forces $v=0$ and (4b) implies $v\geq s$; when $z=1$, (4b) shows $v=s$ and by (4a) we have $v\geq 0$. Therefore, the ReLU activation function is equivalently modeled by the linear constraints in (4) with binary variable $z$. The surrogate model is therefore expressed through these mixed-integer linear constraints:
@@ II-E Overview of the Proposed Framework
12
13Use the input data $(d_{n_{U}1},d_{n_{U}2},\dots,d_{n_{U}N_{P}},w_{n_{W}})$ and the output data $I(u_{n_{U}},w_{n_{W}})$ for $n_{U}=1,2,\dots,N_{U}$ and $n_{W}=1,2,\dots,N_{W}$ to train the MLP with ReLU activation, and obtain the parameters $W_{ij}$ and $B_{ij}$ for $i=1,2,\dots,N_{L}$ and $j=1,2,\dots,N_{Li}$.
@@ III-B Data-Driven Uncertainty Set and Statistical Guarantees
$\mathbb{P}^{N}[\Pr[O\leq O_{x_{0}}\leq O_{\mathcal{U}_{1}}]\geq 1-\varepsilon]\geq 1-\delta.$ (17)
@@ III-C Uncertainty Set Reconstruction
$\mathbb{P}^{N}[\Pr[O\leq O_{x_{1}}\leq O_{\mathcal{U}_{2}}\leq f(x_{0})+\beta]\geq 1-\varepsilon]\geq 1-\delta.$ (19b)
@@ III-D Solution Algorithm
which shows that $\mathcal{U}_{2}$ is polyhedral. Therefore, the two uncertainty sets are polyhedral or ellipsoidal. Problem (11) with these uncertainty sets can be effectively solved by the column-and-constraint generation (C&CG) algorithm [38], which is omitted here for the sake of conciseness.
@@ IV-B1 Benchmark
Based on the performance on the validation dataset, we use the first three components of PCA in the surrogate model for weight optimization. The multilayer perceptron neural network has two hidden layers, and each layer has 16 units. The neural network is trained using the Adam algorithm and the learning rate is 0.001. To help prevent overfitting, we adopt the L2 regularization technique. After about 1000 epochs, the neural network is trained.
The test day for UC is from the test dataset. The MILP problem for weight optimization is solved in 0.02 s and the result is $w=(0.28,0.23,0.49)$. The weight for M3 is the highest, reflecting that the sub-profiles are valuable in power forecasting and robust UC, which is consistent with the results in TABLE I and [17]. With the optimized weight, the UC results are obtained after 121 s, and the optimal value is \$89725. The power outputs of some controllable generators are depicted in Fig. 4, including the pre-dispatch power, the reserve region, and the re-dispatch result in the worst-case scenario. The power in the re-dispatch stage is always within the reserve region.
@@ IV-B2 Comparison
Method Objective (\$) Feasible rate Total cost (\$) Time (s)
@@
As TABLE III shows, SP has the lowest objective value and test total cost. However, its test feasible rate is 8888%, much lower than the desired threshold 9595%, which shows that SP lacks robustness. The other five methods are RO-based and their test feasible rates all meet the requirement. However, RO1 is rather conservative and has the highest objective value and test total cost. RO2 is a traditional data-driven RO method with no statistical guarantees, which shows its theoretical limitation. P1, P2, and the proposed method have statistical guarantees, and the proposed method achieves the lowest optimal objective and test total cost among them. Using uncertainty set reconstruction, the objective decreases 8.308.30%. The integrated forecasting and optimization framework also contributes to improving the objective. The computation time is acceptable for all the methods. Therefore, TABLE III verifies that the proposed method is effective. It outperforms other methods when considering both robustness and optimality. Moreover, the proposed weight optimization and uncertainty set reconstruction processes help improve the performance.
To validate the effectiveness of the surrogate model, we compare the proposed method to a method that directly optimizes the weight. Since the impact of the weight on the evaluated cost of the UC strategy cannot be expressed as a compact function, we choose the PSO algorithm, which is derivative-free and highly flexible [45]. As shown in TABLE III, the PSO algorithm uses a much longer time ($10000$ s) compared to the proposed method ($121$ s), achieving a reduction in the objective value by $0.7\%$ and a decrease in the total cost by $0.1\%$. Thus, the proposed method substantially reduces the computation time to an acceptable duration for UC, with only a mild sacrifice in performance. This underscores the effectiveness of the proposed surrogate model in balancing computational efficiency and performance.
@@
To visualize different uncertainty sets, we examine a specific case involving two random loads. To visualize different uncertainty sets, we consider a special case of two random loads. The uncertainty sets are projected onto a two-dimensional plane. The projections of the uncertainty sets are drawn in Fig. 5. To emphasize the impact of uncertainty set construction and reconstruction, the weight $w$ is fixed to that of C1. The black polygon is the projection of $\mathcal{U}_{0}$. All other uncertainty sets are the intersections of $\mathcal{U}_{0}$ and ellipsoids or polyhedrons. The projected uncertainty set of RO1 is framed by an ellipse that includes all the data points. RO2’s ellipse has the same shape and center but only contains 9595% data points. RO2 does not have a statistical guarantee for its out-of-sample performance. P1 has a larger uncertainty set than RO2 to maintain the statistical guarantees, but P1 does not reconstruct the uncertainty set to decrease the conservativeness. Proposed_1 and Proposed_2 are the projections of the first and second uncertainty sets of the proposed method ($\mathcal{U}_{1}^{\prime}$ and $\mathcal{U}_{2}$ in Algorithm 2), respectively. Proposed_1 is framed by an ellipse similar to that of P1. After the reconstruction, Proposed_2 excludes the right part of the ellipse, thereby omitting some high-cost scenarios in the UC problem. At the same time, Proposed_2 includes additional regions with relatively low costs, ensuring that the data points within it are enough for the statistical guarantees. This approach effectively reduces conservativeness while maintaining statistical guarantees.
@@ IV-C Modified IEEE 118-Bus System
We use a modified IEEE 118-bus system with 54 controllable generators and four wind farms to show the scalability. The parameter settings are $\varepsilon=\delta=$ 5%, $N_{1}=212$, and $N_{2}=124$. More details can be found in [44]. The UC results under different methods are shown in TABLE IV. The method relationships are similar to those in the modified IEEE 30-bus system. In addition, the computation time increases but is still acceptable for day-ahead UC. We also test the computation efficiency under different numbers of random loads in the modified IEEE 118-bus system, as shown in Table V. These results verify the scalability of the proposed method.
@@
Method Objective (\$) Feasible rate Total cost (\$) Time (s)
@@
TABLE V: Computation Efficiency in the Modified IEEE 118-Bus System
@@ Appendix D Proof of Theorem 2
$\mathbb{P}^{N}[\Pr[O\leq O_{x_{1}}\leq O_{\mathcal{U}_{2}}]\geq 1-\varepsilon]\geq 1-\delta.$

### v3

@@ Abstract
The growing uncertainty from renewable power and electricity demand brings significant challenges to unit commitment (UC). While various advanced forecasting and optimization methods have been developed to better predict and address this uncertainty, most previous studies treat forecasting and optimization as separate tasks. This separation can lead to suboptimal results due to misalignment between the objectives of the two tasks. To overcome this challenge, we propose a robust UC framework that integrates forecasting and optimization processes while ensuring statistical guarantees. In the forecasting stage, we combine multiple predictions derived from diverse data sources and methodologies for an improved prediction, aiming to optimize the UC performance. In the optimization stage, the combined prediction is used to construct an uncertainty set with statistical guarantees, based on which the robust UC model is formulated. The optimal robust UC solution provides feedback to refine the weight used for combining multiple predictions. To solve the proposed integrated forecasting-optimization framework efficiently and effectively, we develop a neural network-based surrogate model for acceleration and introduce a reshaping method for the uncertainty set based on the optimization result to reduce conservativeness. Case studies on modified IEEE 30-bus and 118-bus systems demonstrate the advantages of the proposed approach.
@@ I Introduction
If given the probability distribution of uncertainty, stochastic programming (SP) can be applied to UC problems to determine the optimal strategy [1]. However, an accurate probability distribution is difficult to obtain in reality and the inaccuracy of the distribution may jeopardize the feasibility and optimality of SP solutions. Robust optimization (RO) handles this problem by focusing on the worst-case scenario in a pre-defined uncertainty set. Fruitful research has been done in this area: A multi-stage robust UC model was proposed and solved by robust dual dynamic programming in [2]. Polyhedra were used to replace rectangular uncertainty sets to model the uncertainty correlations in [3]. However, RO can be overly conservative when the uncertainty set contains all the possible uncertainty realizations, because extreme scenarios rarely happen but can cause a steep cost increase. The method in [4] estimated quantiles from historical data to establish rectangular uncertainty sets, in which extreme scenarios were excluded. In [5], uncertainty budgets were used to restrict the uncertainty set. A decision-dependent uncertainty set was constructed in [6] to consider the impact of pricing strategies on demand. However, estimated quantiles and uncertainty budgets may still be inaccurate, and the resulting uncertainty set lacks a statistical guarantee. This makes the obtained robust UC strategy less reliable and trustworthy.
@@ II Integrated Forecasting and Optimization Framework
In this section, we first propose an integrated forecasting and optimization framework for a two-stage decision-making problem under uncertainty. In fact, the proposed framework is generally applicable and not restricted to the UC problem. Then in Section III, we introduce the detailed UC model and explain how it fits within the proposed framework.
@@ II-A Forecasting
Consider the forecasting task for a multivariate stochastic process $\mathbf{u}:=(\mathbf{u}_{1},\mathbf{u}_{2},\dots,\mathbf{u}_{t},\dots)$, where $\mathbf{u}_{t}=(u_{it};i\in\mathcal{I})$ is the vector of uncertain variables in period $t$, and $\mathcal{I}$ is the index set. A forecast is then a collection of future values over the horizon of interest, denoted by $\hat{\mathbf{U}}:=(\hat{\mathbf{u}}_{1},\hat{\mathbf{u}}_{2},\dots,\hat{\mathbf{u}}_{T})$. The hat notation indicates these are forecasts, and hence estimates of such future values. Additionally, the corresponding ground truth is denoted by $\mathbf{U}$.
Multiple forecasting methods can be used simultaneously. The prediction by method $m\in\mathcal{M}$ is denoted by $\hat{\mathbf{U}}^{(m)}$, and $\mathcal{M}$ is the set of methods. The final prediction is a convex combination of these predictions, that is, $\hat{\mathbf{U}}(w)=\sum_{m\in\mathcal{M}}w^{(m)}\hat{\mathbf{U}}^{(m)}$, where $w=(w^{(m)};m\in\mathcal{M})$ is the weight vector. It satisfies $w\in\mathcal{W}:=\{w~|~w\geq 0,\sum_{m\in\mathcal{M}}w^{(m)}=1\}$. Using the ground truth $\mathbf{U}$, the forecast error is calculated as $\mathbf{E}(w)=\mathbf{U}-\hat{\mathbf{U}}(w)$, which also depends on the weight $w$. We will denote the collection of $N$ such forecast error groups as $\mathbf{E}^{(1:N)}(w)$. Forecasting methods are not the focus of this paper, and we apply multiple forecasting methods in case studies based on the related literature.
@@ II-D2 MILP-Based Weight Optimization
where $z$ is an auxiliary binary variable and $M>0$ is a sufficiently large constant. When $z=0$, (4a) forces $v=0$ and (4b) implies $v\geq s$; when $z=1$, (4b) shows $v=s$ and by (4a) we have $v\geq 0$. Therefore, the ReLU activation function is equivalently modeled by the linear constraints in (4) with the binary variable $z$. The surrogate model is therefore expressed through these mixed-integer linear constraints:
@@ II-E Overview of the Proposed Framework
12Use the input data $(d_{n_{U}1},d_{n_{U}2},\dots,d_{n_{U}N_{P}},w_{n_{W}})$ and the output data $I(u_{n_{U}},w_{n_{W}})$ for $n_{U}=1,2,\dots,N_{U}$ and $n_{W}=1,2,\dots,N_{W}$ to train the MLP with ReLU activation, and obtain the parameters $W_{ij}$ and $B_{ij}$ for $i=1,2,\dots,N_{L}$ and $j=1,2,\dots,N_{Li}$.
@@ III-B Data-Driven Uncertainty Set and Statistical Guarantees
$\mathbb{P}^{N}[O\leq O_{x_{0}}\leq O_{\mathcal{U}_{1}}]\geq 1-\delta.$ (18)
@@ III-C Uncertainty Set Reconstruction
$\mathbb{P}^{N}[O\leq O_{x_{1}}\leq O_{\mathcal{U}_{2}}\leq f(x_{0})+\beta]\geq 1-\delta.$ (20b)
@@ III-D Solution Algorithm
which shows that $\mathcal{U}_{2}$ is polyhedral. Problem (12) with these uncertainty sets can be effectively solved by the column-and-constraint generation (C&CG) algorithm [38], which is omitted here for the sake of conciseness.
@@ IV-B1 Benchmark
Based on the performance on the validation dataset, we use the first three components of PCA in the surrogate model for weight optimization. The MLP neural network has two hidden layers, and each layer has 16 units. The neural network is trained using the Adam algorithm and the learning rate is 0.001. To help prevent overfitting, we adopt the L2 regularization technique. After about 1000 epochs, the neural network is trained.
The test day for UC is from the test dataset. The MILP problem for weight optimization is solved in 0.02 s and the result is $w=(0.28,0.23,0.49)$. The weight for M3 is the highest, reflecting that the sub-profiles are valuable in power forecasting and robust UC, which is consistent with the results in TABLE I and [17]. With the optimized weight, the UC results are obtained after 121 s, and the optimal value is \$89725. The power outputs of some controllable generators are depicted in Fig. 4, including the pre-dispatch power, the reserve region, and the re-dispatch power in the worst-case scenario. The re-dispatch power is always within the reserve region.
@@ IV-B2 Comparison
Method Objective (\$) Feasible rate Test total cost (\$) Time (s)
@@
As TABLE III shows, SP has the lowest objective value and test total cost. However, its test feasible rate is 8888%, much lower than the desired threshold 9595%, which shows that SP lacks robustness. The other five methods are RO-based and their test feasible rates all meet the requirement. However, RO1 is rather conservative and has the highest objective value and test total cost. RO2 is a traditional data-driven RO method with no statistical guarantees, which shows its theoretical limitation. P1, P2, and the proposed method have statistical guarantees, and the proposed method achieves the lowest optimal objective and test total cost among them. Using uncertainty set reconstruction, the objective decreases 8.308.30%. The integrated forecasting and optimization framework also contributes to improving the objective.
To validate the effectiveness of the surrogate model, we compare the proposed method to a method that directly optimizes the weight. Since the impact of the weight on the evaluated cost of the UC strategy cannot be expressed as a compact function, we choose the PSO algorithm, which is derivative-free and highly flexible [45]. As shown in TABLE III, the PSO algorithm uses a much longer time ($10000$ s) compared to the proposed method ($121$ s), achieving a reduction in the objective value by $0.7\%$ and a decrease in the test total cost by $0.1\%$. Thus, the proposed method substantially reduces the computation time to an acceptable duration for UC, with only a mild sacrifice in performance. This underscores the effectiveness of the proposed surrogate model in balancing computational efficiency and performance.
@@
Therefore, the above method comparison verifies that the proposed method is effective. It outperforms other methods considering the comprehensive performance of robustness, optimality, and computational efficiency. Moreover, the proposed weight optimization and uncertainty set reconstruction processes help improve the performance.
To visualize different uncertainty sets, we examine a specific case involving two random loads. The uncertainty sets are projected onto a two-dimensional plane. The projections of the uncertainty sets are drawn in Fig. 5. To emphasize the impact of uncertainty set construction and reconstruction, the weight $w$ is fixed to that of C1. The black polygon is the projection of $\mathcal{U}_{0}$. All other uncertainty sets are the intersections of $\mathcal{U}_{0}$ and ellipsoids or polyhedrons. The projected uncertainty set of RO1 is framed by an ellipse that includes all the data points. RO2’s ellipse has the same shape and center but only contains 9595% data points. RO2 does not have a statistical guarantee for its out-of-sample performance. P1 has a larger uncertainty set than RO2 to maintain the statistical guarantees, but P1 does not reconstruct the uncertainty set to decrease the conservativeness. Proposed_1 and Proposed_2 are the projections of the first and second uncertainty sets of the proposed method ($\mathcal{U}_{1}^{\prime}$ and $\mathcal{U}_{2}$ in Algorithm 2), respectively. Proposed_1 is framed by an ellipse similar to that of P1. After the reconstruction, Proposed_2 excludes the right part of the ellipse, thereby omitting some high-cost scenarios in the UC problem. At the same time, Proposed_2 includes additional regions with relatively low costs, ensuring that the data points within it are enough for the statistical guarantees. This approach effectively reduces conservativeness while maintaining statistical guarantees.
@@ IV-C Modified IEEE 118-Bus System
We use a modified IEEE 118-bus system with 54 controllable generators and four wind farms to show the scalability. The parameter settings are $\varepsilon=\delta=$ 5%, $N_{1}=212$, and $N_{2}=124$. More details can be found in [44]. The UC results under different methods are shown in TABLE IV. The method relationships are similar to those in the modified IEEE 30-bus system. In addition, the computation time increases but is still acceptable for day-ahead UC. We also test the computational efficiency under different numbers of random loads in the modified IEEE 118-bus system, as shown in Table V. These results verify the scalability of the proposed method.
@@
Method Objective (\$) Feasible rate Test total cost (\$) Time (s)
@@
TABLE V: Computational Efficiency in the Modified IEEE 118-Bus System
@@ Appendix D Proof of Theorem 2
$\mathbb{P}^{N}[O\leq O_{x_{1}}\leq O_{\mathcal{U}_{2}}]\geq 1-\delta.$